the difference of two numbers is 20. find the numbers if one is 3 times the other.

Answer by Maths68(1474)

the difference of two numbers is 20. find the numbers if one is 3 times the other.
Let
Numbers are x and y
x-y=20...........(1)
x=3y.............(2)
Put the value of x in (1)
3y-y=20
2y=20
2y/2=20/2
y=10
Put the value of y in (2)
x=3y
x=3(10)
x=30
Numbers are 30 and 10